OpenStudy (anonymous):

write an equation of a line that is parallel to the given line (8,-5): -5x - 4y=3 HELPP full answer n work plz

OpenStudy (anonymous):

if it is parallel then it must have the same slope, so find the slope of your line by isolating y \[-5x-4y=3\] \[4y=-5x-3\] \[y=-\frac{5}{4}x-\frac{3}{4}\]so you slope is \[-\frac{5}{4}\]

OpenStudy (anonymous):

now use the "point - slope' formula: \[y-y_1=m(x-x_1)\] \[y-(-5)=-\frac{5}{4}(x-8)\] \[y+5=-\frac{5}{4}x+10\] \[y=-\frac{5}{4}x+5\]